# Towards the quark mass dependence of $T_{cc}^{+}$ from lattice QCD

### v1

@@ I Introduction
The past two decades have witnessed the arrival of a wealth of new experimental information on hadronic states with properties at odds with quark model predictions. Such states, conventionally referred to as exotic, are the focus of many theoretical investigations—see, for example, Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9]. For an overview of the experimental situation and theoretical approaches see Ref. [7]. Most of the exotic states that have been discovered contain a heavy quark-antiquark pair ($c\bar{c}$ or $b\bar{b}$), however, experimental signatures of completely new types of exotic states with heavy quarks have recently been detected. This includes the fully charmed tetraquarks seen in the di-charmonium production spectrum [10, 11, 12] and the doubly charmed tetraquark $T_{cc}^{+}$ [13, 14] observed in the proton-proton collisions at the LHC. The latter, being the first representative of a potentially rich family of exotic hadrons with two heavy quarks, has attracted a lot of attention in the hadronic community. A remarkable feature of $T_{cc}^{+}$ is its proximity to the $D^{*+}D^{0}$ and $D^{*0}D^{+}$ thresholds and its quantum numbers $J^{P}=1^{+}$ [13], consistent with $S$-wave scattering in a pseudoscalar-vector mesons system. These two facts taken together suggest that its wave function is dominated by a long-range molecular component [4]. In this scenario, the $T_{cc}^{+}$ peak is associated with a shallow quasibound [^1] We call “quasibound” pole, a pole located below the two-body $DD^{*}$ threshold that would reside on the real axis if $D^{*}$ were stable. We call “binding energy” the difference between the real part of the $T_{cc}^{+}$ pole position in the energy complex plane and the $D^{*+}D^{0}$ threshold while “width” is defined as twice the imaginary part of the pole position. pole with a binding energy of around 350 keV and a width of around 60 keV (see, for example, Refs. [14, 15, 16]). It is instructive to note that, for the physical pion mass, the $T_{cc}^{+}$ width is found to be very sensitive to the 3-body effects related to the pion exchange between $D^{(*)}$ mesons. In particular, the deviation from the above value of the extracted imaginary part of the $T_{cc}^{+}$ pole may come to about 30%, if the 3-body unitarity is violated [15]. Hence, setting up a meaningful low-energy expansion for the $DD^{*}$ scattering amplitude also requires the 3-body effects to be taken into account consistently [17].
@@
Recently, the $T_{cc}^{+}$ state was also studied using the lattice QCD framework in Refs. [18, 19, 20]. In Ref. [18], the HALQCD method was employed in a simulation with $m_{\pi}\simeq 146$ MeV. This procedure involves extracting the $DD^{*}$ scattering potential, which is then used to evaluate the phase shifts above the two-body threshold. In Refs. [19, 20], the conventional Lüscher’s method was employed to extract the $DD^{*}$ phase shifts $\delta(E)$ on lattice QCD ensembles with $m_{\pi}\simeq 280$ and 348 MeV, respectively. All the simulations were performed in the isospin limit and for pion masses exceeding the physical value, such that the $DD\pi$ threshold lies above the $DD^{*}$ threshold and the decay $D^{*}\to D\pi$ is kinematically forbidden. The radiative decay mode $D^{*}\to D\gamma$ is also absent. All the above analyses assumed validity of the effective range expansion (ERE) near the $DD^{*}$ threshold [19, 20, 18]. This rendered a virtual pole [^2] In the given lattice settings with a stable $D^{*}$, this is a generic virtual pole on the real axis below the two-body $DD^{*}$ threshold on the second Riemann sheet with Im$(p)<0$ (here $p$ is the 3-momentum in the $DD^{*}$ system). below threshold in Refs. [19, 18], while the authors of Ref. [20] refrained from extracting the pole position. The validity of the effective range expansion in these unphysically heavy light quark mass set-ups was questioned in Refs. [21, 22], where the authors presented a re-analysis of the finite volume spectrum determined in Ref. [19] without assuming an ERE. In particular, it was pointed out that possible pion exchange interactions between $D^{(*)}$ mesons inter alia imply the presence of the so-called left-hand cuts in the scattering amplitude with the most relevant branch point related to one-pion exchange (see Fig. 1) lying slightly below the $DD^{*}$ threshold, if $m_{\pi}>m_{D^{*}}-m_{D}$.
@@
The present lattice simulation aims at exploring the heavy-quark mass dependence of $DD^{*}$ scattering in the $T_{cc}^{+}$ channel. This has not been studied before. In particular, the simulations in Ref. [20, 18] were performed at the physical charm quark mass and the authors of Ref. [19] considered only two masses very close to the physical point. In this paper, we extract the $DD^{*}$ scattering amplitude for the $cc\bar{u}\bar{d}$ system with $I=0$, $J^{P}=1^{+}$, and $m_{\pi}\simeq 280$ MeV for five heavy-quark masses $m_{c}$, corresponding to $D$-meson masses in the range $m_{D}\simeq 1.7-2.4$ GeV. It is, therefore, an extension of the simulation in Ref. [19] to three additional heavier values of $m_{c}$. We aim to address the hypothesis whether the pole representing $T_{cc}^{+}$ is approaching a bound state pole as the heavy-quark mass is increased, as expected from the theory predictions of a tightly bound $T_{bb}^{-}$ with respect to the $BB^{*}$ threshold—see, for example, the references contained in a recent work [23]. In particular, theory predictions for $T_{bc}$ with $I=0$ and $J^{P}=1^{+}$ do not yet agree on whether the state corresponds to a bound state below the $DB^{*}$ threshold [24, 25, 26] or not [27, 28]. We remark that the reduced mass of the physical $DB^{*}$ system, $m_{r}(DB^{*})\simeq 1.38$ GeV, lies only slightly above the highest reduced mass of the $DD^{*}$ system, $m_{r}(DD^{*})\simeq 1.23$ GeV in our study. Thus the present results may shed light on the nature of the $T_{bc}$, although such conclusions may be influenced by cut-off and other unquantified effects which lie beyond the scope of the present work.
@@
$c_{0}$ [GeV−${}^{-2}$] −$-5.10\pm 0.26$$-4.66\pm 0.21$$-4.42\pm 0.50$$-4.27\pm 0.73$$-5.57\pm 1.92$
$c_{2}$ [GeV−${}^{-4}$] $7.64\pm 0.77$ $6.38\pm 0.63$ $5.57\pm 1.49$ $4.99\pm 2.20$ $7.28\pm 6.03$
[…]
@@
[…]
@@
[…]
@@ II Pion exchange and left-hand cuts
ℒ${\cal L}=\frac{g_{c}}{{2}f_{\pi}}{\bm{D}^{*}}^{\dagger}\cdot{\bm{\nabla}}{\pi^{a}}\tau^{a}D+h.c.\leavevmode\nobreak\ .$ (1)
It renders the one-pion exchange potential
@@ IV Results assuming effective range expansion
In this section, we present the $S$-wave $DD^{*}$ scattering amplitude $T_{0}$ extracted from the eigen-energies above the $DD^{*}$ threshold in the elastic region using Lüscher’s formalism for two-body scattering. The energy dependence is parametrised with an effective range expansion (ERE). Truncating the ERE after the first two terms, we have
@@ IV.1 Heavy-quark mass dependence
Figure 5: Heavy quark mass dependence of the effective range expansion parameters, effective range $r_{0}$ (top left) and inverse scattering length $a_{0}^{-1}$ (bottom left), in units of $E_{\rm th}$. Top right: Correlated dependence of the effective range expansion parameters on the heavy-quark mass. Bottom right: The $DD^{*}$ scattering rate ∝$\propto p|T_{0}|^{2}$ above the threshold in arbitrary units. The left-hand cut branch points are shown as dashed vertical lines and the virtual bound poles are presented as circles.
A simultaneous fit to the levels indicated with dark shaded blue and green circles in Fig. 3 renders the scattering amplitude for ($J=1^{+}$, $l=0$). The momentum-squared or equivalently the energy dependence of $p\cot\delta_{0}$ for the different heavy-quark masses is shown in Fig. 4. For all heavy-quark masses, the quantity $p\cot\delta_{0}$ is increasing with the energy, and the scattering length $a_{0}$ is positive, as evident from panel $\bm{(b)}$ of Fig. 5. The fit results for the ERE parameters are listed in Table 1. Ignoring any left-hand cut effects, the ERE fits cross the unitary parabola $ip$ at the point where
@@ V Effective field theory approach
|$\ket{DD^{*},I=0}=\frac{1}{\sqrt{2}}(\ket{D^{+}D^{*0}}-\ket{D^{0}D^{*+}}).$ (17)
The potential for off-shell one-pion exchange in this system, as depicted in Fig. 1, is provided in Eq. (2), and its $S$-wave projection is given in Eq. (8). To proceed, it is also pertinent to express the pion propagator $1/(u-m_{\pi}^{2})$ in Eq. (8) as a sum of the two contributions in time-ordered perturbation theory (see, for example, Ref. [15]). It gives
@@
$V({\bm{p}},{\bm{p}}^{\prime})=V_{\rm CT}(p,p^{\prime})+V^{S}_{\pi}(p,p^{\prime}).$ (20)
This potential does not have a straightforward representation as a single-argument function in coordinate space and needs to be regularised at short range. In Appendix A we discuss its simplified form to pinpoint some general features relevant for understanding the motion of the $DD^{*}$ scattering amplitude pole. In particular, a typical $V(r)$ extracted from our data has an attractive short-range part and a small barrier at larger distances as sketched in Fig. 8 (see also Fig. 14 below). Note also that, in the total potential in Eq. (20), the short-range part of the pion exchange potential (first term in parentheses in Eq. (7)) can be kept in $V_{\pi}^{S}$ or absorbed by $c_{0}$ in $V_{\rm CT}$; we choose the former option. This ambiguity emphasises the fact that the short-range interaction from the contact terms and pion exchange cannot be disentangled in a scheme-independent way—see Ref. [49] for a detailed discussion.
With the total potential in Eq. (20) at hand we formulate the Lippmann–Schwinger equation for the off-shell $DD^{*}$ scattering amplitude,
@@ VI.1 Heavy-quark mass dependence
The lattice data for the $S$-wave scattering $p\cot\delta_{0}$ for each of the five heavy-quark masses are fitted with the solution of the Lippmann–Schwinger equation in Eq. (21), see Fig. 9. The two fit parameters are the contact terms $c_{0}$ and $c_{2}$ introduced in Eq. (19). Only the three data points lying above the left-hand cut branch point (indicated by the green dashed vertical line in Fig. 9) are included in the fit.888 The fit minimises $\chi^{2}=r_{i}\left(\mathrm{Cov}^{-1}\right)_{ij}r_{j}$, where Cov is the covariance matrix and the residual for the $i$-th level is $r_{i}=(p\cot\delta_{0})^{\rm lat}_{i}-(p\cot\delta_{0})^{\rm EFT}_{i}$.
@@
Having solved the Lippmann–Schwinger equation in Eq. (21), for each heavy-quark mass we obtain a scattering amplitude with the fitted values of the counter terms quoted in Table 1. Next we search for the poles of the on-shell scattering amplitude $T(E)$ defined in Eq. (12) across the complex energy plane $E$ (see Appendix B for technical details). All poles are found to reside on the second Riemann sheet—we list them in Table 1. The pole trajectory with varying heavy-quark mass is visualised in Figs. 11 and 12. In particular,

### v3

@@ I Introduction
The past two decades have witnessed the arrival of a wealth of new experimental information on hadronic states with properties at odds with quark model predictions. Such states, conventionally referred to as exotic, are the focus of many theoretical investigations—see, for example, Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9]. For an overview of the experimental situation and theoretical approaches see Ref. [7]. Most of the exotic states that have been discovered contain a heavy quark-antiquark pair ($c\bar{c}$ or $b\bar{b}$), however, experimental signatures of completely new types of exotic states with heavy quarks have recently been detected. This includes the fully charmed tetraquarks seen in the di-charmonium production spectrum [10, 11, 12] and the doubly charmed tetraquark $T_{cc}^{+}$ [13, 14] observed in the proton-proton collisions at the LHC. The latter, being the first representative of a potentially rich family of exotic hadrons with two heavy quarks, has attracted a lot of attention in the hadronic community. A remarkable feature of $T_{cc}^{+}$ is its proximity to the $D^{*+}D^{0}$ and $D^{*0}D^{+}$ thresholds and its quantum numbers $J^{P}=1^{+}$ [13], consistent with $S$-wave scattering in a pseudoscalar-vector mesons system. These two facts taken together suggest that its wave function is dominated by a long-range molecular component [4]. In this scenario, the $T_{cc}^{+}$ peak is associated with a shallow quasibound [^1] We call “quasibound” pole, a pole located below the two-body $DD^{*}$ threshold that would reside on the real axis if $D^{*}$ were stable. We call “binding energy” the difference between the real part of the $T_{cc}^{+}$ pole position in the energy complex plane and the $D^{*+}D^{0}$ threshold while “width” is defined as twice the imaginary part of the pole position. pole with a binding energy of around 350 keV and a width of around 60 keV (see, for example, Refs. [14, 15, 16]). It is instructive to note that, for the physical pion mass, the $T_{cc}^{+}$ width is found to be very sensitive to the three-body effects related to the pion exchange between $D^{(*)}$ mesons. In particular, the deviation from the above value of the extracted imaginary part of the $T_{cc}^{+}$ pole may come to about 30%, if the three-body unitarity is violated [15]. Hence, setting up a meaningful low-energy expansion for the $DD^{*}$ scattering amplitude also requires the three-body effects to be taken into account consistently [17].
@@
Recently, the $T_{cc}^{+}$ state was also studied using the lattice QCD framework in Refs. [18, 19, 20]. In Ref. [18], the HALQCD method was employed in a simulation with $m_{\pi}\simeq 146$ MeV. This procedure involves extracting the $DD^{*}$ scattering potential, which is then used to evaluate the phase shifts above the two-body threshold. In Refs. [19, 20], the conventional Lüscher’s method was employed to extract the $DD^{*}$ phase shifts $\delta(E)$ on lattice QCD ensembles with $m_{\pi}\simeq 280$ and 348 MeV, respectively. All the simulations were performed in the isospin limit and for pion masses exceeding the physical value, such that the $DD\pi$ threshold lies above the $DD^{*}$ threshold and the decay $D^{*}\to D\pi$ is kinematically forbidden. The radiative decay mode $D^{*}\to D\gamma$ is also absent. All the above analyses assumed validity of the effective range expansion (ERE) near the $DD^{*}$ threshold [19, 20, 18]. This rendered a virtual pole [^2] In the given lattice settings with a stable $D^{*}$, this is a generic virtual pole on the real axis below the two-body $DD^{*}$ threshold on the second Riemann sheet with Im$(p)<0$ (here $p$ is the 3-momentum in the $DD^{*}$ system). below threshold in Refs. [19, 18], while the authors of Ref. [20] refrained from extracting the pole position. The validity of the effective range expansion in these unphysically heavy light quark mass set-ups was questioned in Refs. [21, 22], where the authors presented a re-analysis of the finite volume spectrum determined in Ref. [19] without assuming effective range expansion. In particular, it was pointed out that possible pion exchange interactions between $D^{(*)}$ mesons inter alia imply the presence of the so-called left-hand cuts in the scattering amplitude with the most relevant branch point related to one-pion exchange (see Fig. 1) lying slightly below the $DD^{*}$ threshold, if $m_{\pi}>m_{D^{*}}-m_{D}$.
@@
The present lattice simulation aims at exploring the heavy-quark mass dependence of $DD^{*}$ scattering in the $T_{cc}^{+}$ channel. This has not been studied before. In particular, the simulations in Ref. [20, 18] were performed at the physical charm quark mass and the authors of Ref. [19] considered only two masses very close to the physical point. In this paper, we extract the $DD^{*}$ scattering amplitude for the $cc\bar{u}\bar{d}$ system with $I=0$, $J^{P}=1^{+}$, and $m_{\pi}\simeq 280$ MeV for five heavy-quark masses $m_{c}$, corresponding to $D$-meson masses in the range $m_{D}\simeq 1.7-2.4$ GeV. It is, therefore, an extension of the simulation in Ref. [19] to three additional heavier values of $m_{c}$. We aim to address the hypothesis whether the pole representing $T_{cc}^{+}$ is approaching a bound state pole as the heavy-quark mass is increased, as expected from the theory predictions of a tightly bound $T_{bb}^{-}$ with respect to the $BB^{*}$ threshold—see, for example, the references contained in a recent work [23]. In particular, theory predictions for $T_{bc}$ with $I=0$ and $J^{P}=1^{+}$ do not yet agree on whether the state corresponds to a bound state below the $DB^{*}$ threshold [24, 25, 26] or not [27, 28]. The more recent simulations with either more relevant operators [26] or continuum and chiral extrapolation [25] suggest that a shallow bound state is more likely. We remark that the reduced mass of the physical $DB^{*}$ system, $m_{r}(DB^{*})\simeq 1.38$ GeV, lies only slightly above the highest reduced mass of the $DD^{*}$ system, $m_{r}(DD^{*})\simeq 1.23$ GeV in our study. Thus the present results may shed light on the nature of the $T_{bc}$, although such conclusions may be influenced by cut-off and other unquantified effects which lie beyond the scope of the present work.
@@
$c_{0}$ [GeV-2] −$-5.10\pm 0.26$$-4.66\pm 0.21$$-4.42\pm 0.50$$-4.27\pm 0.73$$-5.57\pm 1.92$
$c_{2}$ [GeV-4] $7.64\pm 0.77$ $6.38\pm 0.63$ $5.57\pm 1.49$ $4.99\pm 2.20$ $7.28\pm 6.03$
pole1: $E^{\rm p}-E_{\rm th}$ [MeV] −$-8.3-9.1\leavevmode\nobreak\ i$$-6.9-8.3\leavevmode\nobreak\ i$$-5.9-7.3\leavevmode\nobreak\ i$$-5.0-6.2\leavevmode\nobreak\ i$$-1.2$
@@
pole2: $E^{\rm p}-E_{\rm th}$ [MeV] −$-8.3+9.0\leavevmode\nobreak\ i$$-6.9+8.3\leavevmode\nobreak\ i$$-5.9+7.3\leavevmode\nobreak\ i$$-5.0+6.2\leavevmode\nobreak\ i$$-5.8$
@@
Table 1: Results for the $S$-wave $DD^{*}$ scattering for five heavy-quark masses at fixed $m_{\pi}\simeq 280$ MeV: Sets 1 and 2 have already been analysed in Ref. [19]; Set 2 is closest to the physical charm quark mass. Top: masses of the scattered mesons $D$ and $D^{*}$ and the employed hopping parameter $\kappa_{c}$, which is related to the bare charm quark mass. Middle: results assuming the effective range expansion. The extrapolation of the effective-range-expansion-based fits to the below-threshold region is questionable due to the presence of a nearby left-hand cut and hence the pole positions should only be taken in the spirit of a consistency check with results from the previous work presented in Ref. [19]. Bottom: improved results based on the effective field theory taking into account the effects of the pion exchange. In the latter case, for Sets 1-4, pole1 and pole2 are complex conjugate poles that correspond to the $T_{cc}^{+}$ as a resonance; for Set 5, the $T_{cc}^{+}$ is associated with the virtual pole1 while pole2 (marked with asterisk) corresponds to another, more remote virtual pole.
@@ II Pion exchange and left-hand cuts
ℒ${\cal L}=\frac{g_{c}}{2f_{\pi}}\left({\bm{D}}^{*\dagger}\cdot{\bm{\nabla}}\pi^{a}\tau^{a}D+D^{\dagger}\tau^{a}{\bm{\nabla}}\pi^{a\dagger}\cdot{\bm{D}}^{*}\right),$ (1)
where
$\tau^{a}\pi^{a}=\begin{pmatrix}\pi^{0}&\sqrt{2}\pi^{+}\\ \sqrt{2}\pi^{-}&-\pi^{0}\end{pmatrix}$
and $f_{\pi}=92.2$ MeV and $g_{c}=0.57$ are the pion decay constant and the $DD^{*}\pi$ coupling, respectively. [^3] Here we quote the values of $f_{\pi}$ and $g_{c}$ relevant for the physical pion mass; the values actually employed in the calculations will be discussed below. Lagrangian (1) renders the one-pion exchange potential
@@ IV Results assuming effective range expansion
In this section, we present the $S$-wave $DD^{*}$ scattering amplitude $T_{0}$ extracted from the eigen-energies above the $DD^{*}$ threshold in the elastic region using Lüscher’s formalism for two-body scattering. The energy dependence is parametrised with an effective range expansion (ERE). Truncating the effective range expansion after the first two terms, we have
@@ IV.1 Heavy-quark mass dependence
Figure 5: Heavy quark mass dependence of the effective range expansion parameters, effective range $r_{0}$ (a) and inverse scattering length $a_{0}^{-1}$ (b), in units of $E_{\rm th}$. (c) Correlated dependence of the effective range expansion parameters on the heavy-quark mass. (d) The $DD^{*}$ scattering rate ∝$\propto p|T_{0}|^{2}$ above the threshold in arbitrary units. The left-hand cut branch points are shown as dashed vertical lines and the virtual bound poles are presented as circles.
A simultaneous fit to the levels indicated with dark shaded blue and green circles in Fig. 3 renders the scattering amplitude for ($J=1^{+}$, $l=0$). The momentum-squared or equivalently the energy dependence of $p\cot\delta_{0}$ for the different heavy-quark masses is shown in Fig. 4. For all heavy-quark masses, the quantity $p\cot\delta_{0}$ is increasing with the energy, and the scattering length $a_{0}$ is positive, as evident from panel $\bm{(b)}$ of Fig. 5. The fit results for the effective range parameters are listed in Table 1. Ignoring any left-hand cut effects, the effective range fits cross the unitary parabola $ip$ at the point where
@@ V Effective field theory approach
|$\ket{DD^{*},I=0}=\frac{1}{\sqrt{2}}(\ket{D^{+}D^{*0}}-\ket{D^{0}D^{*+}}),$ (17)
and the $DD^{*}$ scattering amplitude $T$ in this system is described as illustrated in Fig. 2 and detailed below. In particular, the potential for off-shell one-pion exchange, as depicted in Fig. 1, is provided in Eq. (2), and its $S$-wave projection is given in Eq. (8). To proceed, it is also pertinent to express the pion propagator $1/(u-m_{\pi}^{2})$ in Eq. (8) as a sum of the two contributions in time-ordered perturbation theory (see, for example, Ref. [15]). It gives
@@
$V({\bm{p}},{\bm{p}}^{\prime})=V_{\rm CT}(p,p^{\prime})+V^{S}_{\pi}(p,p^{\prime}),$ (20)
as illustrated in Fig. 2. This potential does not have a straightforward representation as a single-argument function in coordinate space and needs to be regularised at short range. In Appendix A we discuss its simplified form to pinpoint some general features relevant for understanding the motion of the $DD^{*}$ scattering amplitude pole. In particular, a typical $V(r)$ extracted from our data has an attractive short-range part and a small barrier at larger distances as sketched in Fig. 8 (see also Fig. 14 below). Note also that, in the total potential in Eq. (20), the short-range part of the pion exchange potential (first term in parentheses in Eq. (7)) can be kept in $V_{\pi}^{S}$ or absorbed by $c_{0}$ in $V_{\rm CT}$; we choose the former option. This ambiguity emphasises the fact that the short-range interaction from the contact terms and pion exchange cannot be disentangled in a scheme-independent way—see Ref. [49] for a detailed discussion.
With the total potential in Eq. (20) at hand we formulate the Lippmann–Schwinger equation for the off-shell $DD^{*}$ scattering amplitude
@@ VI.1 Heavy-quark mass dependence
The lattice data for the $S$-wave scattering $p\cot\delta_{0}$ for each of the five heavy-quark masses are fitted with the solution of the Lippmann–Schwinger equation in Eq. (21), see Fig. 9. The two fit parameters are the contact terms $c_{0}$ and $c_{2}$ introduced in Eq. (19). Only the three data points lying above the left-hand cut branch point (indicated by the green dashed vertical line in Fig. 9) are included in the fit.999 The fit minimises $\chi^{2}=r_{i}\left(\mathrm{Cov}^{-1}\right)_{ij}r_{j}$, where Cov is the covariance matrix and the residual for the $i$-th level is $r_{i}=(p\cot\delta_{0})^{\rm lat}_{i}-(p\cot\delta_{0})^{\rm EFT}_{i}$.
@@
Having solved the Lippmann–Schwinger equation in Eq. (21), for each heavy-quark mass we obtain a scattering amplitude with the fitted values of the counter terms quoted in Table 1.101010The analysis in Ref. [22] reaffirms the scattering amplitude for our Set 2 also below the left-hand cut. The approach is based on a Hamiltonian formalism employed to extract finite-volume energies in the plane wave basis [53] that is free of the issues related to left-hand cuts. Meanwhile, to stay on a safe ground, in the present analysis, we exclude the lattice points below the left-hand cut until they are independently confirmed using the Lüscher formalism adapted to the presence of the left-hand cuts, for example, as outlined in Refs. [31, 32, 33, 34]. Next we search for the poles of the on-shell scattering amplitude $T(E)$ defined in Eq. (12) across the complex energy plane $E$ (see Appendix B for technical details). All poles are found to reside on the second Riemann sheet—we list them in Table 1. The pole trajectory with varying heavy-quark mass is visualised in Figs. 11 and 12. In particular,